# Making Them a Malicious Database: Exploiting Query Code to Jailbreak Aligned Large Language Models

### v2

@@ Abstract
Making Them a Malicious Database: Exploiting Query Code to Jailbreak Aligned Large Language Models
Qingsong Zou1,3thanks: Equal contribution.  Jingyu Xiao211footnotemark: 1  Qing Li3thanks: Corresponding author.  Zhi Yan4  Yuhang Wang5 Li Xu6  Wenxuan Wang2  Kuofeng Gao1  Ruoyu Li7  Yong Jiang1 1Tsinghua Shenzhen International Graduate School  2The Chinese University of Hong Kong 3Pengcheng Laboratory  4Jilin University  5Southwest University 6University of Electronic Science and Technology of China  7Shenzhen University
@@ 1 Introduction
Specifically, CipherChat (Yuan et al., 2024) uses encryption methods such as the Caesar cipher to translate harmful queries into encrypted text. ArtPrompt (Jiang et al., 2024) replaces sensitive terms with ASCII-style encoding.  Deng et al. (2024) convert sensitive contents into low-resource languages. The essence of these methods lies in inducing the model to generate encrypted outputs, which are then decrypted to harmful text in natural language format. However, they typically require the model to possess knowledge of encryption to understand the prompts or place high requirements on the model’s ability to generate encrypted content. As a result, their attack effectiveness is limited. Therefore, developing an effective and efficient jailbreak attack method remains a critical challenge.
@@
We test QueryAttack on AdvBench (Zou et al., 2023) across both well-known open-source and closed-source LLMs. The experimental results show that QueryAttack effectively bypasses their security defenses, achieving state-of-the-art attack success rates (ASRs). Besides, we provide a visual analysis of QueryAttack’s success and propose a tailored defense method against QueryAttack.
@@ 2 Background
Large language models (LLMs) have demonstrated remarkable generative potential across various fields. However, they are still vulnerable to jailbreak attacks. Jailbreak attacks against LLMs typically involve crafting carefully designed inputs to prompt models to generate and output harmful response, such as instructions that blatantly violate human ethics or the disclosure of sensitive information. Since natural language samples are widely used as safety alignment data during the training phase of LLMs (gpt4), potential malicious users or adversaries can bypass the defense mechanisms of these models by designing prompts based on non-natural language distributions.
@@
Despite the development of numerous defense methods by researchers to mitigate jailbreaking attacks, such as supervised fine-tuning (Wei et al., 2022; Ouyang et al., 2022) and reinforcement learning from human feedback (Sun et al., 2023; Mehrabi et al., 2024). However, recent work has shown that they are not able to completely defend against zero-day jailbreaking attacks (Song et al., 2025; DBLP:conf/coling/WuG0025). Therefore, it is necessary to continue studying the attack vector of LLMs and provide insights for developing new defenses.
@@ 3.1 Query Components Extraction
The query components extraction task can be performed manually by humans or automated using natural language processing tools such as NLTK (Bird et al., 2009). Given the strong text-processing capabilities of LLMs, we choose GPT-4-1106 (OpenAI, 2024a) to perform the component extraction. We design specialized prompts to guide the GPT-4-1106 in components extraction. This prompt is aimed at ensuring the model focuses on the text-processing task rather than the malicious content, and thereby prevents the translation request from being rejected by the model. To achieve this, we employ in-context learning to enable the model to understand the component extraction task and standardize its output. The details of the prompt are provided in the Appendix C.
@@ 3.3 Query Understanding
The purpose of the query learning is to guide the model in understanding the intent behind the query codes and then generate a natural language response. To help the LLMs understand the content, we first establish the context of the conversation. By describing the three query components, we guide the model in building a mapping from the query code to the natural language, and define the conversation within an educational context.
@@
For models with strong understanding of programming languages, we can skip this process and use zero-shot to launch the attack. Finally, we provide detailed guidance in natural language to respond to the query queries. The aim of this process is to have the model answer the relevant knowledge as thoroughly as possible, rather than focusing on understanding and explaining the natural semantics of the prompts. In Appendix A, we provide the complete prompt of this part.
@@ 4.1 Experimental Setup
Victim Models. We test QueryAttack on 13 mainstream large language models: GPT-3.5 (gpt-3.5-turbo) (OpenAI, 2023), GPT-4-1106 (gpt-4-1106-preview) (OpenAI, 2024a), GPT-4o (OpenAI, 2024b), O1 (gpt-o1) (OpenAI, 2024c), Deepseek (deepseek-chat) (DeepSeek-AI et al., 2024), Gemini-flash (gemini-1.5-flash), Gemini-pro (gemini-1.5-pro) (DeepMind, 2024), Llama-3.1-8B (meta-llama-3.1-8B-instruct), Llama-3.1-70B (meta-llama-3.1-70B-instruct), Llama-3.2-1B (meta-llama-3.2-1B-instruct), Llama-3.2-3B (meta-llama-3.2-3B-instruct), Llama-3.2-11B (meta-llama-3.2-11B-vision-instruct) and Llama-3.3-70B (meta-llama-3.3-70B-instruct) (Touvron et al., 2023b; Meta, 2024). To maintain the reproducibility of the results, we set all the temperature to 0.
Datasets. We used AdvBench (Zou et al., 2023) as the dataset for our experiments. This is a harmful behavior dataset that contains 520 different harmful query instances written in natural language. For some experiments, we use a subset of AdvBench which contains 50 representative, non-repetitive harmful instructions refined in (Jiang et al., 2024). We will specify this at the beginning of these parts where the subset is used.
GPT-4-1106 GPT-4o* Llama-3.1-8B* Llama-3.3-70B* Gemini-pro* Gemini-flash*
@@
Harmfulness score (HS). Our evaluation follows the work of (Qi et al., 2024), utilizing the powerful evaluation capabilities of GPT-4 to assign harmfulness scores to the responses generated by the target LLM. The harmfulness score ranges from 1 to 5. To improve the accuracy of GPT-4 Judge, we remove educational statements or disclaimers that may appear in the responses to extract the harmful content, as shown in B.3. The extracted content is then input into GPT-4 Judge for evaluation. Detailed settings can be found in the Appendix B.
@@
Refuse rate (RR). The third evaluation metric is the refuse rate (RR). We follow the setup of (Zou et al., 2023) and use Dict-Judge to assess the number of refusal responses, as shown in Equation 2.
@@ 4.2 Results
QueryAttack achieves SoTA ASR. Table 1 presents the average HS and the ASR of QueryAttack and several baselines on AdvBench (Zou et al., 2023). We demonstrate two configurations of QueryAttack. In the first configuration, denoted as Top 1, QueryAttack uses the programming language style with the highest ASR to construct the query template. In the second configuration, denoted as Ensemble, no restrictions are placed on the programming language styles. From this table, we can observe that, despite specialized safety alignment training in these latest LLMs, QueryAttack successfully bypassed their defenses, inducing responses that violate policies or human values. For GPT-4-1106, the Top 1 configuration achieves a 82.18% ASR and the Ensemble configuration achieves 93.80%. In contrast, the baselines which perform best are only able to bypass GPT-4-1106’s safeguards in up to 81% of cases. The same trend is observed in other models. Except for Llama-3.1-8B, QueryAttack’s ASR is lower than HEA. However, under the Ensemble configuration, the average HS remains higher than that of HEA.
@@ 4.3 Ablation and Analysis
Languages that differ more from natural language is likely to increase QueryAttack’s ASR. Figure 4 shows the average HS and RR obtained by attacking GPT-4-1106, Llama-3.1-70B, Gemini-flash and DeepSeek with templates of different language styles. On GPT-4-1106, Gemini-flash and DeepSeek, different language styles do not show significant variations in average HS. A noticeable decrease is observed when attacking Llama-3.1-70B with URL and SQL-style templates, where a higher RR leads to a lower HS. This may be because these two template styles closely resemble the structure of natural language, making them more likely to trigger existing defenses. Despite this, under the Top 1 configuration, the ASR on Llama-3.1-70B still reachs 76.3%, while the Ensemble configuration achieves 92.9%.
Figure 4: Performance of different language styles.
Gemini-flash GPT-4-1106 GPT-3.5 Llama-3.1-8B
ASR HS ASR HS ASR HS ASR HS
Paraphrase 94% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$ 4.88 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.12}$ 72% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 20\%}$ 4.16 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.76}$ 68% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ 4.56 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.14}$ 90% ${\color[rgb]{1,0,0}\uparrow 2\%}$ 4.86 ${\color[rgb]{1,0,0}\uparrow 0.14}$
Rand-insert 100% ${\color[rgb]{1,0,0}-0\%}$ 5.00 ${\color[rgb]{1,0,0}-0.00}$ 86% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$ 4.76 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.16}$ 66% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 16\%}$ 4.48 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ 72% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ 4.46 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.26}$
Rand-swap 100% ${\color[rgb]{1,0,0}-0\%}$ 5.00 ${\color[rgb]{1,0,0}-0.00}$ 94% ${\color[rgb]{1,0,0}\uparrow 2\%}$ 4.88 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.04}$ 54% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 28\%}$ 4.12 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.58}$ 70% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 16\%}$ 4.50 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$
Rand-patch 100% ${\color[rgb]{1,0,0}-0\%}$ 5.00 ${\color[rgb]{1,0,0}-0.00}$ 94% ${\color[rgb]{1,0,0}\uparrow 2\%}$ 4.90 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.02}$ 64% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 18\%}$ 4.48 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ 72% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ 4.50 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$
Ours 36% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 74\%}$ 3.56 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.44}$ 28% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 64\%}$ 3.10 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.82}$ 76% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$ 4.68 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.02}$ 34% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 52\%}$ 3.38 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.34}$
Table 2: QueryAttack’s average ASR / HS against defense baselines. The differences between the defense and no defense are indicated by arrows.

### v3

# Qingsong Zou1,2  Jingyu Xiao311footnotemark: 1  Qing Li2  Zhi Yan4  Yuhang Wang5 Li Xu6  Wenxuan Wang3  Kuofeng Gao1  Ruoyu Li7  Yong Jiang1, [^2] footnotemark: 2 1Tsinghua Shenzhen International Graduate School  2Pengcheng Laboratory

@@ Abstract
[Uncaptioned image] QueryAttack: Jailbreaking Aligned Large Language Models Using Structured Non-natural Query Language
Qingsong Zou1,2thanks: Equal contribution.  Jingyu Xiao311footnotemark: 1  Qing Li2thanks: These authors contributed equally to this work as co-corresponding authors.  Zhi Yan4  Yuhang Wang5 Li Xu6  Wenxuan Wang3  Kuofeng Gao1  Ruoyu Li7  Yong Jiang1, [^2] footnotemark: 2 1Tsinghua Shenzhen International Graduate School  2Pengcheng Laboratory 3The Chinese University of Hong Kong  4Jilin University  5Southwest University 6University of Electronic Science and Technology of China  7Shenzhen University zouqs21@mails.tsinghua.edu.cn  jyxiao@link.cuhk.edu.hk  liq@pcl.ac.cn  jiangy@sz.tsinghua.edu.cn
@@ 1 Introduction
Specifically, CipherChat (Yuan et al., 2024) uses encryption methods such as the Caesar cipher to translate harmful queries into encrypted text. ArtPrompt (Jiang et al., 2024) replaces sensitive terms with ASCII-style encoding.  (Deng et al., 2024) convert sensitive contents into low-resource languages. The essence of these methods lies in inducing the model to generate encrypted outputs, which are then decrypted to harmful text in natural language format. However, they typically require the model to possess knowledge of encryption to understand the prompts or place high requirements on the model’s ability to generate encrypted content. As a result, their attack effectiveness is limited. To illustrate this point, we design a simple yet clear experiment to test whether some mainstream large language models can effectively understand and generate encrypted text, as shown in Appendix A. The results show that some models may fail to achieve both objectives simultaneously and cause jailbreak failure. Therefore, developing an effective and efficient jailbreak attack method remains a critical challenge.
@@
We test QueryAttack on AdvBench (Zou et al., 2023b) across both well-known open-source and closed-source LLMs. The experimental results show that QueryAttack effectively bypasses their security defenses, achieving state-of-the-art attack success rates (ASRs). Besides, we provide a visual analysis of QueryAttack’s success and propose a tailored defense method against QueryAttack.
@@ 2 Background
Large language models (LLMs) have demonstrated remarkable generative potential across various fields. However, they are still vulnerable to jailbreak attacks. Jailbreak attacks against LLMs typically involve crafting carefully designed inputs to prompt models to generate and output harmful response, such as instructions that blatantly violate human ethics or the disclosure of sensitive information. Since natural language samples are widely used as safety alignment data during the training phase of LLMs (OpenAI, 2024a), potential malicious users or adversaries can bypass the defense mechanisms of these models by designing prompts based on non-natural language distributions.
@@
Despite the development of numerous defense methods by researchers to mitigate jailbreaking attacks, such as supervised fine-tuning (Wei et al., 2022; Ouyang et al., 2022) and reinforcement learning from human feedback (Sun et al., 2023; Mehrabi et al., 2024). However, recent work has shown that they are not able to completely defend against zero-day jailbreaking attacks (Song et al., 2025; Wu et al., 2025). Therefore, it is necessary to continue studying the attack vector of LLMs and provide insights for developing new defenses.
@@ 3.1 Query Components Extraction
The query components extraction task can be performed manually by humans or automated using natural language processing tools such as NLTK (Bird et al., 2009). Given the strong text-processing capabilities of LLMs, we choose GPT-4-1106 (OpenAI, 2024a) to perform the component extraction. We design specialized prompts to guide the GPT-4-1106 in components extraction. This prompt is aimed at ensuring the model focuses on the text-processing task rather than the malicious content, and thereby prevents the translation request from being rejected by the model. To achieve this, we employ in-context learning to enable the model to understand the component extraction task and standardize its output. The details of the prompt are provided in the Appendix D.
@@ 3.3 Query Understanding
In-context learning is also applied to this step. However, unlike in query components extraction, the purpose of the query learning is to guide the model in understanding the intent behind the query codes and then generate a natural language response. To help the LLMs understand the content, we first establish the context of the conversation. By describing the three query components, we guide the model in building a mapping from the query code to the natural language, and define the conversation within an educational context.
@@
For models with strong understanding of programming languages, we can skip this process and use zero-shot to launch the attack. Finally, we provide detailed guidance in natural language to respond to the query queries. The aim of this process is to have the model answer the relevant knowledge as thoroughly as possible, rather than focusing on understanding and explaining the natural semantics of the prompts. In Appendix B, we provide the complete prompt of this part.
@@ 4.1 Experimental Setup
Victim Models. We test QueryAttack on 14 mainstream large language models: GPT-3.5 (gpt-3.5-turbo) (OpenAI, 2023), GPT-4-1106 (gpt-4-1106-preview) (OpenAI, 2024a), GPT-4o (OpenAI, 2024b), O1 (gpt-o1) (OpenAI, 2024c), DeepSeek (deepseek-chat) (DeepSeek-AI et al., 2024), DeepSeek-R1 (DeepSeek-R1) (DeepSeek-AI et al., 2025), Gemini-flash (gemini-1.5-flash), Gemini-pro (gemini-1.5-pro) (DeepMind, 2024), Llama-3.1-8B (meta-llama-3.1-8B-instruct), Llama-3.1-70B (meta-llama-3.1-70B-instruct), Llama-3.2-1B (meta-llama-3.2-1B-instruct), Llama-3.2-3B (meta-llama-3.2-3B-instruct), Llama-3.2-11B (meta-llama-3.2-11B-vision-instruct) and Llama-3.3-70B (meta-llama-3.3-70B-instruct) (Touvron et al., 2023b; Meta, 2024). To maintain the reproducibility of the results, we set all the temperature to 0.
Datasets. We use AdvBench (Zou et al., 2023b) and HEx-PHI Zou et al. (2023a) as the dataset for our experiments. AdvBench is a harmful behavior dataset that contains 520 different harmful query instances written in natural language. For HEx-PHI, we use the subset of it refined in (Jiang et al., 2024) for evaluation. We use QueryAttack to convert these harmful queries written in natural language into structured non-query language to attack the target LLMs. For some experiments, we use a subset of AdvBench which contains 50 representative, non-repetitive harmful instructions refined in (Jiang et al., 2024). We will specify this at the beginning of these parts where the subset is used.
Method GPT-4-1106 GPT-4o* Llama-3.1-8B* Llama-3.3-70B* Gemini-pro* Gemini-flash*
@@
Harmfulness score (HS). Our evaluation follows the work of (Qi et al., 2024), utilizing the powerful evaluation capabilities of GPT-4 to assign harmfulness scores to the responses generated by the target LLM. The harmfulness score ranges from 1 to 5. To improve the accuracy of GPT-4 Judge, we remove educational statements or disclaimers that may appear in the responses to extract the harmful content, as shown in C.3. The extracted content is then input into GPT-4 Judge for evaluation. Detailed settings can be found in the Appendix C.
@@
Refuse rate (RR). The third evaluation metric is the refuse rate (RR). We follow the setup of (Zou et al., 2023b) and use Dict-Judge to assess the number of refusal responses, as shown in Equation 2.
@@ 4.2 Results
QueryAttack achieves SoTA ASR. Table 1 presents the average HS and the ASR of QueryAttack and several baselines on AdvBench (Zou et al., 2023b). We demonstrate two configurations of QueryAttack. In the first configuration, denoted as Top 1, QueryAttack uses the programming language style with the highest ASR to construct the query template. In the second configuration, denoted as Ensemble, no restrictions are placed on the programming language styles. From this table, we can observe that, despite specialized safety alignment training in these latest LLMs, QueryAttack successfully bypassed their defenses, inducing responses that violate policies or human values. For GPT-4-1106, the Top 1 configuration achieves a 82.18% ASR and the Ensemble configuration achieves 93.80%. In contrast, the baselines which perform best are only able to bypass GPT-4-1106’s safeguards in up to 81% of cases. The same trend is observed in other models. Except for Llama-3.1-8B, QueryAttack’s ASR is lower than HEA. However, under the Ensemble configuration, the average HS remains higher than that of HEA.
We conduct additional experiments using another widely-used public dataset, HEx-PHI (Zou et al., 2023a). For cost considerations, we employ the HEx-PHI subset defined in ArtPrompt (Jiang et al., 2024) (containing 11 categories with 10 samples each, totaling 110 samples) to test the harm score / attack success rate across four mainstream LLMs using Python, C++, and SQL, as shown in the Table 2. In most cases, QueryAttack still achieves high ASR, demonstrating its effectiveness on different datasets.
C++ Python SQL Ensemble
Gemini-1.5-flash 4.76 / 92.73% 4.74 / 89.09% 4.33 / 73.64% 4.84 / 94.55%
GPT-4-1106 4.60 / 80.00% 4.38 / 72.73% 4.32 / 76.36% 4.65 / 83.64%
DeepSeek-R1 4.81 / 88.18% 4.72 / 88.18% 4.71 / 87.27% 4.92 / 93.64%
Llama-3.1-70B 4.38 / 76.37% 3.98 / 64.55% 2.09 / 14.55% 4.51 / 82.73%
Table 2: QueryAttack’s attack performance on HEx-PHI (Zou et al., 2023a).
@@ 4.3 Ablation and Analysis
Languages that differ more from natural language is likely to increase QueryAttack’s ASR. Figure 4 shows the average HS and RR obtained by attacking GPT-4-1106, Llama-3.1-70B, Gemini-flash and DeepSeek with templates of different language styles. On GPT-4-1106, Gemini-flash and DeepSeek, different language styles do not show significant variations in average HS. A noticeable decrease is observed when attacking Llama-3.1-70B with URL and SQL-style templates, where a higher RR leads to a lower HS. This may be because these two template styles closely resemble the structure of natural language, making them more likely to trigger existing defenses. Despite this, under the Top 1 configuration, the ASR on Llama-3.1-70B still reaches 76.3%, while the Ensemble configuration achieves 92.9%.
@@
Gemini-flash GPT-4-1106 GPT-3.5 Llama-3.1-8B
ASR HS ASR HS ASR HS ASR HS
Paraphrase $94\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$ $4.88_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.12}$ $72\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 20\%}$ $4.16_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.76}$ $68\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ $4.56_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.14}$ $90\%_{\color[rgb]{1,0,0}\uparrow 2\%}$ $4.86_{\color[rgb]{1,0,0}\uparrow 0.14}$
Rand-insert $100\%_{\color[rgb]{1,0,0}-0\%}$ $5.00_{\color[rgb]{1,0,0}-0.00}$ $86\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$ $4.76_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.16}$ $66\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 16\%}$ $4.48_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ $72\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ $4.46_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.26}$
Rand-swap $100\%_{\color[rgb]{1,0,0}-0\%}$ $5.00_{\color[rgb]{1,0,0}-0.00}$ $94\%_{\color[rgb]{1,0,0}\uparrow 2\%}$ $4.88_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.04}$ $54\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 28\%}$ $4.12_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.58}$ $70\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 16\%}$ $4.50_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$
Rand-patch $100\%_{\color[rgb]{1,0,0}-0\%}$ $5.00_{\color[rgb]{1,0,0}-0.00}$ $94\%_{\color[rgb]{1,0,0}\uparrow 2\%}$ $4.90_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.02}$ $64\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 18\%}$ $4.48_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ $72\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ $4.50_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$
Ours $36\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 74\%}$ $3.56_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.44}$ $28\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 64\%}$ $3.10_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.82}$ $76\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$ $4.68_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.02}$ $34\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 52\%}$ $3.38_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.34}$
Table 4: QueryAttack’s average ASR / HS against defense baselines. The differences between the defense and no defense are indicated by arrows. Shaded ellipses represents the 95% confidence interval (CI) for each data distribution.
Embedding Differences Between Natural Language and QueryAttack. In addition to the analysis of attention mechanism, we further discuss the embedding differences between structured non-natural query languages and natural language in LLMs. Specifically, we use the embedding layer of Llama-3.2-1B (Meta, 2024) and OpenAI’s text-embedding-3-large model (Neelakantan et al., 2022) to embed both structured non-natural queries and their corresponding natural language queries. We randomly sampled 50 samples from AdvBench (Zou et al., 2023b) and compared the embedding differences between the original natural language queries and structured non-natural queries based on Java and C, and then analyze and visualize their differences using t-SNE, with the results presented in Figure 6.
The results demonstrate a significant divergence in embeddings between structured non-natural query languages and natural language in LLMs, suggesting that QueryAttack could bypass deployed safety alignment mechanisms with a high probability.
(a) Llama-3.2-1B (C and Java)
(b) OpenAI-Emb-Lg (C and Java)
Figure 6: The t-SNE visualizations of the embedding representations of original queries and QueryAttack using Llama-3.2-1B and text-embedding-3-large.